# Testing Hypotheses regarding Covariance and Correlation matrices with the R package CovCorTest

### v1

@@ Introduction
To calculate the required quantiles to use this test statistic, a simulation-based method similar to the Taylor-based Monte Carlo approach is employed to mimic the limit distribution. To this aim for both groups $i=1,2$ independent realisations $\boldsymbol{\Lambda}_{i}\sim\mathcal{N}_{p}({\bf 0}_{p},\widehat{\boldsymbol{\Sigma}}_{i})$ are generated, and subsequently transformed to $\boldsymbol{\Lambda}_{i}^{Tay}$. The extensive formula for the used transformation and their derivation is also part of the online supplement in Sattler and Pauly (2023). From this, we compute $\boldsymbol{T}^{1,Tay}=\sqrt{N}(n_{1}^{-1/2}\boldsymbol{\Lambda}_{1}^{Tay}-n_{2}^{-1/2}\boldsymbol{\Lambda}_{2}^{Tay})$ and repeat this $B\in{\mathbb{N}}$ times, allowing to derive empirical $\beta$ quantiles for each component, denoted as $q_{\ell,\beta}$. Subsequent, the local $\beta$ has to be determined in a way, that it controls the family-wise type-I-error rate by $\alpha$, which is assured through the approach of Munko et al. (2024) by using
${\widetilde{\beta}=\max\left(\beta\in\mathcal{B}\Big{\lvert}\sum\limits_{b=1}^{B}\max\limits_{\ell=1,...,p}\frac{\left(1-11\left(q_{\ell,\frac{\beta}{2}}\leq T_{\ell}^{b,Tay}\leq q_{\ell,1-\frac{\beta}{2}}\right)\right)}{B}\leq\alpha\right)}$
with =$\mathcal{B}=\left\{0,\frac{1}{B},...,\frac{B-1}{B}\right\}$.
@@
$\max\left[\max\limits_{\ell=1,...,d}\left(1-11\left(q_{\ell,\frac{\widetilde{\beta}}{2}}\leq T_{\ell}\leq q_{\ell,1-\frac{\widetilde{\beta}}{2}}\right)\right),\max\limits_{\ell=d+1,...,p}\left(1-11\left(q_{\ell,\frac{\widetilde{\beta}}{2}}\leq T_{\ell}\leq q_{\ell,1-\frac{\widetilde{\beta}}{2}}\right)\right)\right],$
@@
      test_covariance_structure and test_correlation_structure allow to choose from a list of predefined hypotheses, which depends on the considered test respectively function:
@@
      Correlation Structure: The predefined structures of the correlation matrices are heterogenous autoregressive ("hautoregressive" or "har"), heterogenous toeplitz ("htoeplitz" or "htoep"), heterogenous compound symmetry ("hcompoundsymmetry", "hcs") and diagonal ("diagonal" or "diag").
    • For test_covariance and test_correlation, more experienced users could specify custom hypothesis through a hypothesis matrix together with a corresponding vector, as introduced in Sattler et al. (2022) or Sattler and Pauly (2023), especially for less standard hypotheses which increase flexibility. Both refer to the corresponding kind of upper triangular vectorization, and unlike in other procedures, the hypothesis matrix here has neither to be a projection matrix nor square.
@@
  • repetitions: The number of Monte Carlo or bootstrap repetitions used in the approximation of the quantiles. This number substantially influences both, the test’s performance as well as its computation time. Since a small number of repetitions increases the influence of the randomness on the results, a predefined value of 1000 is a good trade-off and substantially smaller numbers are not recommended.
@@
Here we must differentiate between the combined test, and all other implemented test functions.
For the latter ones, the output consists of different parts, which are mostly general and contains more descriptive information about the conducted test, like the investigated null hypothesis (through the label from the list or the hypothesis matrix $\boldsymbol{C}$ and the corresponding vector $\boldsymbol{\zeta}$), the used method for calculating the critical values, the used number of repetitions and the vector containing the group sample sizes. Moreover, the estimated covariance matrix for the test statistic, as well as the value of the test statistic and the p-value, are returned. The print function only gives a short and compact summary of the conducted test, containing the number of groups, the tested hypothesis, the resampling method together with the used number of repetitions, and especially the value of the test statistic and the p-value.
For the combined test, the hypothesis, the number of groups, and the resampling method are fixed, since only one hypothesis can be tested for two groups. Moreover, based on the used Taylor-based Monte-Carlo approach, no covariance estimator of the test statistic is used and, therefore, also not returned. Moreover, the test statistic here is vector-valued and non-standardised, making the components hardly comparable. Since conclusions based on the value of the test statistic are usually not possible here, these values are not printed to avoid misinterpretations. However, this test results not only in one p-value but also in three of them. A \$pvalue_Variances for the local hypothesis of equal variances, a \$pvalue_Correlations for the local hypothesis of equal correlations, and \$pvalue_Total for the global hypothesis of equal covariance matrices (and therefore both, equal correlations and equal variances). All of them, are returned and printed through the function.
@@
Due to the sample sizes, we will check whether there is homogeneity of covariance matrices for women with these diagnoses, leading to $a=3$ unbalanced groups. The test is done two times, using the Monte-Carlo approach and the parametric bootstrap as a comparison, while for the latter, both kinds of specifying the hypothesis of interest are used.
@@
As expected, since the value of the test statistic is not affected through the resampling method, both values coincide, while we get slightly different p-values. With p-values of 0.024 resp 0.023, the null hypothesis of equal covariance matrices in all groups would be rejected to the usual level of $5\%$.
@@
After comparing the covariance matrices for women with these diagnoses, the correlation matrices are also investigated for equality. In addition to the parametric bootstrap, the test is also conducted for the Taylor-based Monte-Carlo approach, specifically developed for testing vectorized correlation matrices.
@@
Instead of testing the equality of covariances matrices and the equality of correlation matrices for two groups separately, the combined test from Sattler and Pauly (2023) allows us to do both at the same time. This allows for the null hypothesis of equal covariance matrices, to conduct whether the difference lies in the correlations and, therefore, the dependency, the variances, or even both.
@@
From the resulting three p-values, we see that both covariance matrices differ, while only a difference in the correlation can be conducted, and it was not possible to prove a difference in the variances.
@@
In this analysis, we want to investigate whether the position of the measuring points influences the measured values. If we reject the hypothesis of the covariance matrix being a compound symmetry matrix, it is verified that there is an effect of the position. Moreover, to check whether also the correlations between the measure points are equal, we additionally check for a heterogeneous compound symmetry structure by using the correlation matrix.
@@
Since both structures are rejected to a usual significance level of $5\%$, not only an influence of the measure points’ location was verified, but also the correlation between them differs.
@@
In the presented paper, we illustrate the functionality and scope of the R package CovCorTest, which provides a comprehensive framework for hypothesis testing concerning covariance and correlation matrices in multivariate data analysis. It contains tests for both parameters and different hypotheses in a uniform way, which excels through its applicability to large class of possible hypotheses and less distribution assumptions. Apart from multi-group settings, single groups are also allowed as well as factorial designs.
In the main, it contains functions for five different tests: regarding covariance matrices, correlation matrices, a combined test for equality of covariance and correlation matrices, and tests regarding their corresponding structures. To this aim, quantiles based, among other things, on resampling techniques and Monte-Carlo approaches are used. Although the Monte-Carlo approach leads to the least time-consuming procedure, a parametric bootstrap for covariances and a Taylor-based Monte-Carlo approach for correlations are preferable for smaller sample sizes.
To include actual developments and procedures in this area, we aim to update this package on a regular basis. Planned future developments include further hypotheses (especially additional structures and patterns) and new test statistics, such as multiple contrast tests for hypotheses regarding covariance and correlation matrices. Also, the combined correlation and covariance test could be expanded to accommodate more than two groups, and tests for multiple structures at once, as mentioned in Sattler and Dobler (2024), could be incorporated. Further, the output of a confidence ellipsoid is planned as a further extension, together with the corresponding plot. Also, a graphical user interface (GUI) is possible to simplify the usage, especially for more applied users. Finally, the results from Sattler and Rosenbaum (2025) could be realized in the future, to reduce calculation effort. This, would allow a higher number of bootstrap and Monte-Carlo repetitions and thereby increase the performance for large datasets.

### v2

@@ Introduction
To calculate the required quantiles to use this test statistic, a simulation-based method similar to the Taylor-based Monte Carlo approach is employed to mimic the limit distribution. To this aim for both groups $i=1,2$ independent realisations $\boldsymbol{\Lambda}_{i}\sim\mathcal{N}_{p}({\bf 0}_{p},\widehat{\boldsymbol{\Sigma}}_{i})$ are generated, and subsequently transformed to $\boldsymbol{\Lambda}_{i}^{Tay}$. The extensive formula for the used transformation and its derivation is also part of the online supplement in Sattler and Pauly (2023). From this, we compute $\boldsymbol{T}^{1,Tay}=\sqrt{N}(n_{1}^{-1/2}\boldsymbol{\Lambda}_{1}^{Tay}-n_{2}^{-1/2}\boldsymbol{\Lambda}_{2}^{Tay})$ and repeat this $B\in{\mathbb{N}}$ times, allowing to derive empirical $\beta$ quantiles for each component, denoted as $q_{\ell,\beta}$. Subsequently, the local $\beta$ has to be determined in a way that it controls the family-wise type-I-error rate by $\alpha$, which is assured through the approach of Munko et al. (2024) by using
$\widetilde{\beta}=\max\left(\beta\in\mathcal{B}\Big\lvert\sum\limits_{b=1}^{B}\max\limits_{\ell=1,...,p}\frac{\left(1-1\hskip-3.01385pt1\left(q_{\ell,\frac{\beta}{2}}\leq T_{\ell}^{b,Tay}\leq q_{\ell,1-\frac{\beta}{2}}\right)\right)}{B}\leq\alpha\right)$}
with $\mathcal{B}=\left\{0,\frac{1}{B},...,\frac{B-1}{B}\right\}$.
@@
$\max\left[\max\limits_{\ell=1,...,d}\left(1-1\hskip-3.01385pt1\left(q_{\ell,\frac{\widetilde{\beta}}{2}}\leq T_{\ell}\leq q_{\ell,1-\frac{\widetilde{\beta}}{2}}\right)\right),\max\limits_{\ell=d+1,...,p}\left(1-1\hskip-3.01385pt1\left(q_{\ell,\frac{\widetilde{\beta}}{2}}\leq T_{\ell}\leq q_{\ell,1-\frac{\widetilde{\beta}}{2}}\right)\right)\right],$
@@
      test_covariance_structure and test_correlation_structure allow to choose from a list of predefined hypotheses, which depends on the considered test, respectively function:
@@
      Correlation Structure: The predefined structures of the correlation matrices are heterogeneous autoregressive ("hautoregressive" or "har"), heterogeneous toeplitz ("htoeplitz" or "htoep"), heterogeneous compound symmetry ("hcompoundsymmetry", "hcs") and diagonal ("diagonal" or "diag").
    • For test_covariance and test_correlation, more experienced users could specify custom hypotheses through a hypothesis matrix together with a corresponding vector, as introduced in Sattler et al. (2022) or Sattler and Pauly (2023), especially for less standard hypotheses, which increase flexibility. Both refer to the corresponding kind of upper triangular vectorisation, and unlike in other procedures, the hypothesis matrix here has neither to be a projection matrix nor square. This especially allows structures from the linear covariance structure model and the linear correlation structure model by using the function get_hypothesis.
@@
  • repetitions: The number of Monte Carlo or bootstrap repetitions used in the approximation of the quantiles. This number substantially influences both the test’s performance as well as its computation time. Since a small number of repetitions increases the influence of the randomness on the results, a predefined value of 1000 is a good trade-off and substantially smaller numbers are not recommended.
@@
Here, we must differentiate between the combined test and all other implemented test functions.
For the latter ones, the output consists of different parts, which are mostly general and contains more descriptive information about the conducted test, like the investigated null hypothesis (through the label from the list or the hypothesis matrix $\boldsymbol{C}$ and the corresponding vector $\boldsymbol{\zeta}$), the used method for calculating the critical values, the used number of repetitions and the vector containing the group sample sizes. Moreover, the estimated covariance matrix for the test statistic, as well as the value of the test statistic and the p-value, are returned. The print function only gives a short and compact summary of the conducted test, containing the number of groups, the tested hypothesis, the resampling method together with the used number of repetitions, and especially the value of the test statistic and p-value.
For the combined test, the hypothesis, the number of groups, and the resampling method are fixed, since only one hypothesis can be tested for two groups. Moreover, based on the used Taylor-based Monte-Carlo approach, no covariance estimator of the test statistic is used and, therefore, also not returned. Moreover, the test statistic here is vector-valued and non-standardised, making the components hardly comparable. Since conclusions based on the value of the test statistic are usually not possible here, these values are not printed to avoid misinterpretations. However, this test results in not only one p-value but also in three of them. A \$pvalue_Variances for the local hypothesis of equal variances, a \$pvalue_Correlations for the local hypothesis of equal correlations, and \$pvalue_Total for the global hypothesis of equal covariance matrices (and therefore both, equal correlations and equal variances). All of them are returned and printed through the function.
@@
Due to the sample sizes, we will check whether there is homogeneity of covariance matrices for women with these diagnoses, leading to $a=3$ unbalanced groups. The test is performed twice, using the Monte-Carlo approach and the parametric bootstrap as a comparison. For the latter, both kinds of specifying the hypothesis of interest are used.
@@
As expected, since the value of the test statistic is not affected by the resampling method, both values coincide, while we get slightly different p-values. With p-values of 0.024 resp 0.023, the null hypothesis of equal covariance matrices in all groups would be rejected at the usual level of $5\%$.
@@
After comparing the covariance matrices for women with these diagnoses, the correlation matrices are also investigated for equality. In addition to the parametric bootstrap, the test is also conducted for the Taylor-based Monte-Carlo approach, specifically developed for testing vectorised correlation matrices.
@@
Instead of testing the equality of covariance matrices and the equality of correlation matrices for two groups separately, the combined test from Sattler and Pauly (2023) allows us to do both at the same time. This allows for the null hypothesis of equal covariance matrices to determine whether the difference lies in the correlations and, therefore, the dependency, the variances, or even both.
@@
From the resulting three p-values, we see that both covariance matrices differ, while only a difference in the correlation can be determined, and it was not possible to prove a difference in the variances.
@@
In this analysis, we want to investigate whether the position of the measuring points influences the measured values. If we reject the hypothesis of the covariance matrix being a compound symmetry matrix, it is verified that there is an effect of the position. Moreover, to check whether the correlations between the measure points are also equal, we additionally check for a heterogeneous compound symmetry structure by using the correlation matrix.
@@
Since both structures are rejected at a usual significance level of $5\%$, not only was the influence of the measure points’ location verified, but also that the correlation between them differs.
We can also test the compound symmetry structure of the covariance matrix, using a hypothesis formulation based on get_hypothesis and TestCovariance.
> X <- t(as.matrix(EEGwide[EEGwide\$sex == "W" & EEGwide\$diagnosis == "AD",
         "complexity_temporal","complexity_frontal", "complexity_central")]))
> v0=rep(0,21)
> vauxilariy=c(1,rep(0,5),1,rep(0,4),1,rep(0,3),1,rep(0,2),1,0,1)
> V=cbind(vauxilariy,1-vauxilariy)
> Hypothesis=get_hypothesis(v0,V)
> set.seed(123)
> test_covariance(X = X,C = Hypothesis\$hypothesis_matrix, Xi = Hypothesis\$hypothesis_vector,
                  method = "MC", repetitions = 1000)
         Covariance Test
            one group
 Hypothesis:            C * v = Xi
 Teststatistic value:   4.0926
 p-value:               p = 0.006
Although this is the same test procedure, data set, and seed, the p-value differs. This indicates that for the ATS, the test decision typically depends on the representation of the null hypothesis, as also described in Sattler and Zimmermann (2023).
@@
In the presented paper, we illustrate the functionality and scope of the R package CovCorTest, which provides a comprehensive framework for hypothesis testing concerning covariance and correlation matrices in multivariate data analysis. It contains tests for both parameters and different hypotheses in a uniform way, which excels through its applicability to a large class of possible hypotheses and fewer distribution assumptions. Apart from multi-group settings, single groups are also allowed, as well as factorial designs. In the main, it contains functions for five different tests: regarding covariance matrices, correlation matrices, a combined test for equality of covariance and correlation matrices, and tests regarding their corresponding structures. To this end, quantiles, based on various techniques including resampling and Monte Carlo approaches, are used. Although the Monte-Carlo approach leads to the least time-consuming procedure, a parametric bootstrap for covariances and a Taylor-based Monte-Carlo approach for correlations are preferable for smaller sample sizes.
To include actual developments and procedures in this area, we aim to update this package on a regular basis. Planned future developments include further hypotheses (especially additional structures and patterns) and new test statistics, such as multiple contrast tests for hypotheses regarding covariance and correlation matrices. Also, the combined correlation and covariance test could be expanded to accommodate more than two groups, and tests for multiple structures at once, as mentioned in Sattler and Dobler (2024), could be incorporated. Further, the output of a confidence ellipsoid is planned as a further extension, together with the corresponding plot. Also, a graphical user interface (GUI) is possible to simplify the usage, especially for more applied users. Finally, the results from Sattler and Rosenbaum (2025) could be realised in the future, to reduce calculation effort. This would allow a higher number of bootstrap and Monte-Carlo repetitions and thereby increase the performance for large datasets.